# Multiple Object Detection and Tracking in Panoramic Videos for Cycling Safety Analysis

### v2

@@ 4.1.2 Object Detection with Pre-trained Models
For the object detection model, this study uses YOLO12 (Tian et al., 2025a), which has state-of-the-art object detection accuracy at the time of writing. Tian et al. (2025b) provides several models with different structures designed for different speed and accuracy requirements. In this step, YOLO12n pre-trained on the COCO dataset was chosen since it is light-weight but still performs well on the validation dataset.
@@ 4.1.3 Reprojection of the Detection
As Figure 4(a) shows, to map all the bounding boxes, which are at least 15° away from the edges back to the original image, every point on the borders of these boxes was reprojected with the approach above. However, this process distorts the edges of the boxes, which makes it difficult to output and evaluate the predictions. Therefore, minimum bounding rectangles are used represent the boxes.
Due to the boundary continuity issue, detections crossing the centre lines of the sub-images projected along 180° are shown twice near the left and the right edges, respectively, in equirectangular views. Thus, when dealing with such detections, as illustrated in Figure 4(b), instead of directly projecting the whole box, this study first divided each of the boxes into two parts along the centre line, and then projected them separately, so that the detection outputs can be displayed and evaluated correctly.
@@ 4.1.4 Merging the Reprojected Bounding Boxes
In the final step, to merge bounding boxes of objects that cannot be fully displayed in a single sub-image, the boxes with at least one side coinciding with a boundary of a sub-image were extracted. According to Figure 4(c), first, those boxes that are entirely within the 30-degree-wide overlapped areas were removed, as they must be included in another bounding box with smaller distortions. Then, the boxes of the same category are combined, which are tangential to consecutive boundaries, into their MBRs. At the same time, the score of each new bounding box was calculated as the area-weighted average.
@@ 4.2 Improving Multiple Object Tracking on Panoramic Cycling Videos
The second step of the proposed methodology is to improve the applicability of StrongSORT (Du et al., 2023), one of the most recently proposed MOT models following the tracking-by-detection paradigm, to panoramic cycling videos. As shown in Figure 5, the following modifications were done: First, improved Yolo12 object detection model was set as the detector; Then, support for category information and boundary continuity were introduced into the process of data association, according to the characteristics of panoramic videos.
@@ 4.2.1 StrongSORT
According to Du et al. (2023), tracks in StrongSORT have two states: ‘confirmed’ and ‘unconfirmed’ — only if a track gets matched in a certain number of consecutive frames will its state be switched to confirmed. For each frame in a video, StrongSORT sends $n$ detections and $m$ existing confirmed tracks into a vanilla global linear assignment module. The module extracts appearance features from each detection and updates the appearance state in an exponential moving average (EMA) manner proposed in Wang et al. (2020b), instead of storing them with feature bank mechanism as in DeepSORT (Wojke et al., 2017). The framework tracks with a pre-trained ReID model and calculates the appearance feature distance and Mahalanobis distance between them, whose weighted averages are then used to form an $m\times n$ cost matrix. Following this, the Hungarian algorithm is applied to associate the tracks and detection according to the values in the cost matrix, and the matched tracks are then updated by NSA Kalman filters (Du et al., 2021) using the information of the corresponding detection.
@@ 4.2.2 Introducing the Support for Category Information
As mentioned above, when associating the detections with tracks, StrongSORT ignores the category information of the detection, treating objects of different categories equally. Thus, tracks and detections can be wrongly matched when their corresponding values in the cost matrices are small. If this unmodified model is applied to panoramic cycling videos; such wrong associations can be seen when objects of different categories alternate at a short distance. To address this issue, category information is included into the tracker. In this way, every track is initialized with the category label of the detection. When computing cost matrices for matching, category filtering was performed. This involved setting values representing the distances between tracks and new detections with different or non-adjacent categories to an arbitrarily large number. This ensures that the existing tracks are only matched with the detections of the same or adjacent categories. Classes of objects like cars and trucks, bicycles and motorcycles are considered as adjacent categories. Sometimes the detector may assign an object its adjacent category in consecutive frames and then a track with incorrect category will be generated. By relaxing the condition of penalisation to adjacent categories, objects that are associated with an incorrect track can be re-matched with true category in later frames.
@@ 4.2.3 Introducing the Support for Boundary Continuity
To track the detections which are divided into two parts (i.e., objects crossing the centre lines of the sub-images projected along 180°), this research proposes an approach referring to the annotation method for panoramic MOT designed in subsection 3.2: the first step is to move each of the left bounding boxes to the right and merge it with the right box; Then, input the combined bounding boxes into StrongSORT directly, despite the fact they are beyond the boundaries. In addition, to avoid ID switching caused by objects that leave from one side of the image and return from the other side, the distance calculation process of StrongSORT was also improved. As Figure 6 shows, for calculating the Mahalanobis or IOU distance between a predicted track and a detection, the proposed method first duplicates the detection on two virtual images extended from both sides of the original image; Then, it calculates the distances from the track to these three bounding boxes respectively and selects the shortest one to fill the cost matrix.
Figure 6: Preparations for the improved distance calculation process (suitable for Mahalanobis distance and IOU distance)
@@ 4.3.1 Movement Direction Recognition
According to the definition of equirectangular projection (Snyder, 1997), objects behind the camera are shown on the left and right sides in a panoramic video. In contrast, objects in the front are shown in the middle. Therefore, if an object’s current position is closer to the centre line than its position in the last frame, it is moving forward relative to the camera. Similarly, if an object moves away from the centre, it is moving backward.
@@ 4.3.2 Overtaking Behaviour Detection
The project defined overtaking behaviour as the act of a motor vehicle moving forwards and completely passing the cyclist. In an equirectangular video, this definition can be abstracted as the occasion in which the front and rear of a vehicle on the left (or right) side of the image cross the -90° (or 90°) longitude line in order.
Figure 7: Workflow of the application which can detect the overtaking behaviour of the surrounding vehicles
Figure 8: Flow chart of the overtaking behaviour detection for a specific object
Figure 8 shows how the application was designed to realise the overtaking behaviour detection on a specific object in the panoramic videos: Take a vehicle on the left side of the video as an example, if the right side of its bounding box crosses the -90° longitude line for the first time as it is moving forwards, it will be assigned with an unconfirmed (ongoing) overtake. The overtake will remain unconfirmed until the left boundary of the bounding box passes the -90° line or the right border goes back to the left of the line, which means the overtake is finished (confirmed) or failed, respectively.
@@ 5.1 Evaluation of the Model for Improving Object Detection on Panoramic Cycling Images
The COCO-style annotated ground truth is used to assess the performance of the pre-trained YOLO12n before and after the method was applied to the dataset described in Section 3 under different input resolution settings (640, 1280, 1920). Mean average precision (mAPs) and average precision (APs) metrics of different sizes were recorded and compared, as higher APs indicate better object detection performance. In addition, to assess the effect of merging the bounding boxes, which is an innovation compared to the method in Yang et al. (2018)’s study, the performance of models with and without combining the bounding boxes is also compared.
@@
Table 1: Average Precision Metrics of YOLO12n before and after the Proposed Methods are Applied. Bolded values represent the best results
@@ 5.2 Evaluation of the Model for Improving Multiple Object Tracking on Panoramic Cycling Videos
As shown in Table 1, YOLO12n achieves high accuracy with moderate memory usage when the input size is 1280. Therefore, it was used as the detector of StrongSORT in the following analyses.
Then, using the improved YOLO as detectors, models before and after the implementations of support for category information and boundary continuity are compared, whose results are shown in Table 2.
@@
Original YOLO + Original StrongSORT 51.5% 76.1% 38.9% 818 12285 101 304 43.7% 0.201
Improved YOLO + Original StrongSORT 45.4% 70.4% 33.6% 1030 13265 141 470 38.5% 0.191
Improved YOLO + Original StrongSORT (Boundary Support) 46.0% 72.5% 33.8% 766 13272 134 471 39.6% 0.191
Improved YOLO + StrongSORT (Category Support) 45.5% 70.5% 33.6% 1021 13266 142 470 38.5% 0.191
Improved YOLO + StrongSORT (Category and Boundary Support) 50.0 % 78.8% 36.6% 544 13100 90 373 41.5% 0.190
Table 2: Evaluation Metrics of StrongSORT before and after the Supports for Category Information and Boundary Continuity are Introduced. Best results in bold
As shown in the table, incorporating category information reduces false positives. Consequently, IDF1 improves marginally from 45.4% to 45.5%, with the number of ID switches and MOTA remaining nearly unchanged. Although the improving effect on metrics seems relatively slight, observing the output videos of the models with and without the support for category information shows that some incorrect data matching has been successfully avoided. On the other hand, the inclusion of boundary support can significantly reduce the number of false positives by almost 25.6 per cent. The IDF1 and MOTA are also be improved by 0.6% and 0.9%, respectively.
@@
These enhancements are also evident in the output tracking videos: With the support for boundary continuity, StrongSORT can now treat the two parts of the same object as a whole; In addition, using the improved distance calculation method, StrongSORT is now capable of retaining the IDs of the objects that leave from one side of the image and return from the other side.
@@ 5.3 Evaluation of the Automated Overtaking Behaviour Detection
In this experiment, we applied our improved YOLO12n ++ StrongSORT(Category and Boundary Support) framework to an overtaking detection task. Here, only cars, trucks and buses were selected as overtaking vehicles due to their more frequent involvement in crashes involving cyclists. The improved algorithm was tested on one hundred 15-second $360^{\circ}$ video clips collected under different riding conditions in London, UK, as mentioned in previous section 3. The input resolution for YOLO12n is 1280. Table 3 presents the test results as a confusion matrix.
Predicted Positive Predicted Negative
Actual Positive 40 10
Actual Negative 8 42
Table 3: Evaluation Results of Overtaking Behaviour Detection - Experiment 2
It can be observed that the algorithm achieves a precision of
$\text{Precision}=\frac{TP}{TP+FP}=\frac{40}{40+8}=0.83,$
and a recall (sensitivity) of
$\text{Recall}=\frac{TP}{TP+FN}=\frac{40}{40+10}=0.80.$
The F-score is computed as
$F\text{-score}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\times 0.83\times 0.80}{0.83+0.80}=0.81.$
@@ 6.1 Key Contributions
Although Table 1 indicates that the performance of the proposed method with input resolution 640 can also be achieved by the original models with larger input resolution, once the input resolution is as large as 1280 or 1920, the effect can not be replaced anymore. In other words, it increases the upper limit of accuracy that an object detection model can achieve on a panoramic image. Conversely, since the performance of the proposed object detection model on small and medium objects is always similar to (or even better than) that of the original models with higher input resolution, for the same accuracy requirements, the proposed model consumes fewer computing resources. Thus, it can achieve competitive accuracy even on low-end GPUs (Yang et al., 2018).
The proposed tracking method for panoramic videos produces the annotation and results output of the objects which go across the boundaries in an equirectangular image without violating the MOT16 annotation rules (Milan et al., 2016). In this way, the existing MOT metrics can still be used for evaluating the tracking results on panoramic videos. Moreover, based on the performance improvement brought by the proposed object detection models, the introduction of category information and support for boundary continuity further enhances the applicability of StrongSORT to panoramic videos by reducing the number of false positives and ID switches.
@@ 6.2 Limitations and Future Work
Similar to its predecessor DeepSORT, StrongSORT also uses a deep appearance ReID network trained on pedestrian datasets (Wojke et al., 2017; Du et al., 2023). Although Zuraimi and Zaman (2021)’s research has demonstrated that it is possible to use a pretrained ReID descriptor on pedestrian images to extract appearance features on other objects such as vehicles, the results presented here show that when the angle of a vehicle with respect to the camera changes, some ID switching may occur. As Figure 11 shows, using the current pretrained ReID network, the vanilla matching module in StrongSORT fails in matching the front and rear of these two cars. Referring to Li et al. (2021), this research plans to construct a vehicle image dataset and re-train a ReID network with it, which may improve the performance of state-of-the-art tracking algorithms like StrongSORT in associating the vehicles according to their appearance.
Figure 11: Examples of ID switching when the angle of a vehicle toward the camera changes
@@ 7 Conclusion
In cycling safety, most existing studies applied computer vision models to ‘traditional’ videos with a limited field of view for automated road user behaviour analysis. To enable analysis of the complete environment surrounding the rider, this research demonstrates an end‑to‑end approach for analysing cyclist safety by combining object detection and MOT in panoramic equirectangular videos. Existing pre-trained object detection models are applied to four perspective sub-images projected from equirectangular images. When evaluated on the annotated panoramic dataset, the performance of YOLO12n is improved under any input resolution setting. For MOT, a method that improved YOLO is used as the detector and support for boundary continuity and category information is implemented in StrongSORT, reducing false positives by 33.5% and id switches by 10.9%. The tracking results of the improved StrongSORT are implemented in the case of automated overtaking detection, which achieves high recall and precision on the testing videos.
@@
Although the proposed methodology has been applied to overtaking detection, the framework can detect and track objects defined as classes in a pre-trained object detection model, such as pedestrians, bicycles, buses and motorcycles. Therefore, in the policy perspectives, the method can quantify compliance with minimum passing‑distance policies and identify contexts (time of day, vehicle class) with elevated risk—e.g., challenges detecting black cars at night and bounding‑box instability for buses highlight specific night‑time and large‑vehicle contexts where targeted enforcement and operator training may be most beneficial. Moreover, fine‑grained evidence on where passes initiate and complete relative to the rider can inform design details (lane widths, buffers, separation types), signal timings, and bus‑stop bypasses. The boundary‑aware tracking particularly helps follow vehicles across the full 360°, supporting evidence on mid‑block and junction approaches. With lightweight deployment, tracking outputs can back real‑time rider alerts—e.g., early warnings of a vehicle preparing to overtake or a “door‑zone” advisory—alongside post‑ride feedback that helps cyclists adjust lane position or route choice. The paper outlines this real‑time potential and the need for efficient models on edge devices .
@@
Overall, agencies and practitioners can use the proposed 360° detection‑tracking‑overtaking pipeline to generate near‑miss surrogates from routine rides, map and monitor hazardous passing, evaluate infrastructure and policy changes when crashes are rare, and support both targeted enforcement and rider-facing warning systems—accelerating evidence‑led cycling safety decisions. However, two technical issues need to be highlighted as the current limitations: (i) detection of large objects occasionally degrades because projections split them across sub‑images. Increasing overlap or sub‑image count may help, balanced against runtime, and (ii) night‑time/black vehicles and buses can challenge the overtaking rule set due to missed detections and unstable boxes.

### v3

@@ 4.1.2 Object Detection with Pre-trained Models
For the object detection model, this study uses YOLO12tian2025yolov12, which has state-of-the-art object detection accuracy at the time of writing. In this step, YOLO12n pre-trained on the COCO dataset was chosen since it is light-weight but still performs well on the validation dataset.
@@ 4.1.3 Reprojection of the Detection
As Figure 3a shows, to map all the bounding boxes, which are at least 15 away from the edges back to the original image, every point on the borders of these boxes was reprojected with the approach above. However, this process distorts the edges of the boxes, which makes it difficult to output and evaluate the predictions. Therefore, minimum bounding rectangles are used represent the boxes.
Due to the boundary continuity issue, detections crossing the centre lines of the sub-images projected along 180 are shown twice near the left and the right edges, respectively, in equirectangular views. Thus, when dealing with such detections, as illustrated in Figure 3b, instead of directly projecting the whole box, this study first divided each of the boxes into two parts along the centre line, and then projected them separately, so that the detection outputs can be displayed and evaluated correctly.
@@ 4.1.4 Merging the Reprojected Bounding Boxes
In the final step, to merge bounding boxes of objects that cannot be fully displayed in a single sub-image, the boxes with at least one side coinciding with a boundary of a sub-image were extracted. According to Figure 3c, first, those boxes that are entirely within the 30-degree-wide overlapped areas were removed, as they must be included in another bounding box with smaller distortions. Then, the boxes of the same category are combined, which are tangential to consecutive boundaries. At the same time, the score of each new bounding box was calculated as the area-weighted average.
@@ 4.2 Improving Multiple Object Tracking on Panoramic Cycling Videos
The second step of the proposed methodology is to improve the applicability of StrongSORT du2023strongsort, one of the most recently proposed MOT models following the tracking-by-detection paradigm, to panoramic cycling videos. As shown in Figure 4, the following modifications were done: First, improved Yolo12 object detection model was set as the detector; Then, support for category information and boundary continuity were introduced into the process of data association, according to the characteristics of panoramic videos.
@@ 4.2.1 StrongSORT
According to du2023strongsort, tracks in StrongSORT have two states: ‘confirmed’ and ‘unconfirmed’ — only if a track gets matched in a certain number of consecutive frames will its state be switched to confirmed. For each frame in a video, StrongSORT sends $n$ detections and $m$ existing confirmed tracks into a vanilla global linear assignment module. The module extracts appearance features from each detection and updates the appearance state in an exponential moving average (EMA) manner proposed in wang2020towards, instead of storing them with feature bank mechanism as in DeepSORT wojke2017simple. The framework tracks with a pre-trained ReID model and calculates the appearance feature distance and Mahalanobis distance between them, whose weighted averages are then used to form an $m\times n$ cost matrix. Following this, the Hungarian algorithm is applied to associate the tracks and detection according to the values in the cost matrix, and the matched tracks are then updated by NSA Kalman filters du2021giaotracker using the information of the corresponding detection.
@@
Figure 4: Workflow of improved object tracking model using StrongSORT as backbone for better tracking performance on panoramic cycling videos.
@@ 4.2.2 Introducing the Support for Category Information
As mentioned above, when associating the detections with tracks, StrongSORT ignores the category information of the detection, treating objects of different categories equally. Thus, tracks and detections can be wrongly matched when their corresponding values in the cost matrices are small. If this unmodified model is applied to panoramic cycling videos; such wrong associations can be seen when objects of different categories alternate at a short distance.
To address this issue, category information is included into the tracker. In this way, every track is initialized with the category label of the detection. When computing cost matrices for matching, category filtering was performed. This involved setting values representing the distances between tracks and new detections with different or non-adjacent categories to an arbitrarily large number. This ensures that the existing tracks are only matched with the detections of the same or adjacent categories. Classes of objects like cars and trucks, bicycles and motorcycles are considered as adjacent categories. Sometimes the detector may assign an object its adjacent category in consecutive frames and then a track with incorrect category will be generated. By relaxing the condition of penalisation to adjacent categories, objects that are associated with an incorrect track can be re-matched with true category in later frames.
@@ 4.2.3 Introducing the Support for Boundary Continuity
To track the detections which are divided into two parts (i.e., objects crossing the centre lines of the sub-images projected along 180), this research proposes an approach referring to the annotation method for panoramic MOT designed in subsection 3.2: the first step is to move each of the left bounding boxes to the right and merge it with the right box; Then, input the combined bounding boxes into StrongSORT directly, despite the fact they are beyond the boundaries. In addition, to avoid ID switching caused by objects that leave from one side of the image and return from the other side, the distance calculation process of StrongSORT was also improved. As Figure 5 shows, for calculating the Mahalanobis or IOU distance between a predicted track and a detection, the proposed method first duplicates the detection on two virtual images extended from both sides of the original image; Then, it calculates the distances from the track to these three bounding boxes respectively and selects the shortest one to fill the cost matrix.
Figure 5: Preparations for the improved distance calculation process (suitable for Mahalanobis distance and IOU distance).
@@ 4.3.1 Movement Direction Recognition
According to the definition of equirectangular projection snyder1997flattening, objects behind the camera are shown on the left and right sides in a panoramic video. In contrast, objects in the front are shown in the middle. Therefore, if an object’s current position is closer to the centre line than its position in the last frame, it is moving forward relative to the camera. Similarly, if an object moves away from the centre, it is moving backward.
@@ 4.3.2 Overtaking Behaviour Detection
The project defined overtaking behaviour as the act of a motor vehicle moving forwards and completely passing the cyclist. In an equirectangular video, this definition can be abstracted as the occasion in which the front and rear of a vehicle on the left (or right) side of the image cross the -90 (or 90) longitude line in order.
Figure 6: Workflow of the application which can detect the overtaking behaviour of the surrounding vehicles.
Figure 7: Flow chart of the overtaking behaviour detection for a specific object.
Figure 7 shows how the application was designed to realise the overtaking behaviour detection on a specific object in the panoramic videos: Take a vehicle on the left side of the video as an example, if the right side of its bounding box crosses the -90 longitude line for the first time as it is moving forwards, it will be assigned with an unconfirmed (ongoing) overtake. The overtake will remain unconfirmed until the left boundary of the bounding box passes the -90 line or the right border goes back to the left of the line, which means the overtake is finished (confirmed) or failed, respectively.
@@ 5.1 Evaluation of the Model for Improving Object Detection on Panoramic Cycling Images
The COCO-style annotated ground truth is used to assess the performance of the pre-trained YOLO12n before and after the method was applied to the dataset described in Section 3 under different input resolution settings (640, 1280, 1920). Mean average precision (mAPs) and average precision (APs) metrics of different sizes were recorded and compared, as higher APs indicate better object detection performance. In addition, to assess the effect of merging the bounding boxes, which is an innovation compared to the method in yang2018object’s study, the performance of models with and without combining the bounding boxes is also compared.
@@
Table 1: Average Precision Metrics of YOLO12n before and after the Proposed Methods are Applied. Bolded values represent the best results.
@@
{tablenotes}
@@ 5.2 Evaluation of the Model for Improving Multiple Object Tracking on Panoramic Cycling Videos
As shown in Table 1, YOLO12n achieves high accuracy with moderate memory usage when the input size is 1280. Therefore, it was used as the detector of StrongSORT in the following analyses. Using the improved YOLO as detectors, models before and after the implementations of support for category information and boundary continuity are compared, whose results are shown in Table 2.
Table 2: Evaluation Metrics of StrongSORT before and after the Supports for Category Information and Boundary Continuity are introduced.
@@
Original YOLO + StrongSORT 51.5% 76.1% 38.9% 818 12285 101 304 43.7% 0.201
Improved YOLO + StrongSORT 45.4% 70.4% 33.6% 1030 13265 141 470 38.5% 0.191
[…]
Improved YOLO + StrongSORT w/ SB and SC 50.0% 78.8% 36.6% 544 13100 90 373 41.5% 0.190
{tablenotes}
Support for Boundary Continuity
Support for Category Information
As shown in the table 2, incorporating category information reduces false positives. Consequently, IDF1 improves marginally from 45.4% to 45.5%, with the number of ID switches and MOTA remaining nearly unchanged. Although the improving effect on metrics seems relatively slight, observing the output videos of the models with and without the support for category information shows that some incorrect data matching has been successfully avoided. On the other hand, the inclusion of boundary support can significantly reduce the number of false positives by almost 25.6 per cent. The IDF1 and MOTA are also be improved by 0.6% and 0.9%, respectively.
@@ 5.3 Evaluation of the Automated Overtaking Behaviour Detection
In this experiment, we applied our improved YOLO12n ++ StrongSORT framework to overtaking behaviour detection. Here, only cars, trucks and buses were selected as overtaking vehicles due to their more frequent involvement in crashes involving cyclists. The improved algorithm was tested on one hundred 15-second $360^{\circ}$ video clips collected under different riding conditions in London, UK, as mentioned in previous section 3. The input resolution for YOLO12n is 1280. Table 3 presents the test results as a confusion matrix.
Table 3: Performance of overtaking behaviour detection using proposed detection framework powered by YOLO + StrongSORT.
[…]
Improved YOLO + StrongSORT 79.0% 80.9% 76.0% 78.4% 38 41 9 12
[…]
Improved YOLO + StrongSORT w/ SB and SC 82.0% 83.3% 80.0% 81.6% 40 42 8 10
{tablenotes}
Support for Boundary Continuity
Support for Category Information
It can be observed that Boundary Support and Category Support for the tracking model (in our case StrongSORT) achieves the best performance for the downstream overtaking detection task with a precision of $\text{Precision}=TP/(TP+FP)=40/(40+8)=0.83,$ and a recall (sensitivity) of $\text{Recall}=TP/(TP+FN)=40/(40+10)=0.80$. The F-score is computed as $F\text{-score}=2\times\text{Precision}\times\text{Recall}/(\text{Precision}+\text{Recall})=2\times 0.83\times 0.80/(0.83+0.80)=0.81$. Removing SB and SC results in a modest but consistent decline in overtaking detection performance, with accuracy dropping from 82.0% to 79.0%, precision from 83.3% to 80.9%, recall from 80.0% to 76.0%, and F1-score from 81.6% to 78.4%. Meanwhile, as the qualitative study shown in Figure 8, the overtaking detection can demonstrate a decent performance on detecting positive overtaking events.
Figure 8: This figure illustrates five representative examples of automatically detected overtaking behaviours from the cyclist’s egocentric perspective. Each row corresponds to a single overtaking event and consists of three consecutive frames, in which the overtaking vehicle is highlighted by a red bounding box.
@@ 6.1 Key Contributions
Although Table 1 indicates that the performance of the proposed method with input resolution 640 can also be achieved by the original models with larger input resolution, once the input resolution is as large as 1280 or 1920, the effect can not be replaced anymore. In other words, it increases the upper limit of accuracy that an object detection model can achieve on a panoramic image. Conversely, since the performance of the proposed object detection model on small and medium objects is always similar to (or even better than) that of the original models with higher input resolution, for the same accuracy requirements, the proposed model consumes fewer computing resources. Thus, it can achieve competitive accuracy even on low-end GPUs yang2018object.
The proposed tracking method for panoramic videos produces the annotation and results output of the objects which go across the boundaries in an equirectangular image without violating the MOT16 annotation rules milan2016mot16benchmarkmultiobjecttracking. In this way, the existing MOT metrics can still be used for evaluating the tracking results on panoramic videos. Moreover, based on the performance improvement brought by the proposed object detection models, the introduction of category information and support for boundary continuity further enhances the applicability of StrongSORT to panoramic videos by reducing the number of false positives and ID switches.
@@ 6.2 Limitations and Future Work
Similar to its predecessor DeepSORT, StrongSORT also uses a deep appearance ReID network trained on pedestrian datasets wojke2017simple, du2023strongsort. Although zuraimi2021vehicle’s research has demonstrated that it is possible to use a pretrained ReID descriptor on pedestrian images to extract appearance features on other objects such as vehicles, the results presented here show that when the angle of a vehicle with respect to the camera changes, some ID switching may occur. Referring to Li2021, this research plans to construct a vehicle image dataset and re-train a ReID network with it, which may improve the performance of state-of-the-art tracking algorithms like StrongSORT in associating the vehicles according to their appearance.
@@ 7 Conclusion
In cycling safety, most existing studies applied computer vision models to ‘traditional’ videos with a limited field of view for automated road user behaviour analysis. To enable analysis of the complete environment surrounding the rider, this research demonstrates an end-to-end approach for analysing cyclist safety by combining object detection and MOT in panoramic equirectangular videos. Existing pre-trained object detection models are applied to four perspective sub-images projected from equirectangular images. When evaluated on the annotated panoramic dataset, the performance of YOLO12n is improved under any input resolution setting. For MOT, a method that improved YOLO is used as the detector and support for boundary continuity and category information is implemented in StrongSORT, reducing false positives by 33.5% and id switches by 10.9%. The tracking results of the improved StrongSORT are implemented in the case of automated overtaking detection, which achieves high recall and precision on the testing videos.
@@
Although the proposed methodology has been applied to overtaking detection, the framework can detect and track objects defined as classes in a pre-trained object detection model, such as pedestrians, bicycles, buses and motorcycles. Therefore, in the policy perspectives, the method can quantify compliance with minimum passing-distance policies and identify contexts (time of day, vehicle class) with elevated risk — e.g., challenges detecting black cars at night and bounding-box instability for buses highlight specific night-time and large-vehicle contexts where targeted enforcement and operator training may be most beneficial. Moreover, fine-grained evidence on where passes initiate and complete relative to the rider can inform design details (lane widths, buffers, separation types), signal timings, and bus-stop bypasses. The boundary-aware tracking particularly helps follow vehicles across the full 360, supporting evidence on mid-block and junction approaches. With lightweight deployment, tracking outputs can back real-time rider alerts — e.g., early warnings of a vehicle preparing to overtake or a “door-zone” advisory — alongside post-ride feedback that helps cyclists adjust lane position or route choice. The paper outlines this real-time potential and the need for efficient models on edge devices .
@@
Overall, agencies and practitioners can use the proposed 360 detection-tracking-overtaking pipeline to generate near-miss surrogates from routine rides, map and monitor hazardous passing, evaluate infrastructure and policy changes when crashes are rare, and support both targeted enforcement and rider-facing warning systems — accelerating evidence-led cycling safety decisions. However, two technical issues need to be highlighted as the current limitations: (i) detection of large objects occasionally degrades because projections split them across sub-images. Increasing overlap or sub-image count may help, balanced against runtime, and (ii) night-time/black vehicles and buses can challenge the overtaking rule set due to missed detections and unstable boxes.
[…]
*Data availability statement The video data from the 100 Cyclists Project used in this study are not publicly available to protect the privacy of the participants. A sample video collected by the research team is available on the project’s Github page at https://github.com/SpaceTimeLab/360_object_tracking, along with the code used to generate the results.
[…]
*Acknowledgments The authors would like to thank the participants of the 100 Cyclists Project for volunteering their time and contributing valuable data to this research. They would also like to thank the reviewers for their feedback, which have helped to improve the manuscript.
[…]
*Conflict of interest
The authors declare no potential conflict of interests.